# Differentially private projection-depth-based medians

### v1

@@ Appendix A Algorithm
Appendix B Proofs from Section 3.1
Prior to proving the main results, we first establish probability bounds that will be used throughout the following proofs.
@@
Let $\hat{\alpha}_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}-\inf_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}$ and $\alpha_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\mu_{u}-\inf_{u\in\mathbb{S}^{d-1}}\mu_{u}$.
@@
    $\Pr\Big{(}\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}\sigma(\tilde{\nu})\leq\frac{c_{1}}{2}\Big{)}\lesssim a_{1,d}e^{-a_{2,d,k}nc_{1}/4}+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4}).$ (B.1)
@@
    $\Pr\left(\widehat{a}_{\mu}>4c_{3}\right)\lesssim\zeta_{n,d}(c_{3}).$ (B.2)
@@
    $\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\geq 2c_{2}\right)\lesssim\zeta_{n,d}(c_{2}).$ (B.3)
@@
    $\Pr\left(\inf_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\leq c_{1}/2\right)\lesssim\zeta_{n,d}(\frac{c_{1}}{2}).$ (B.4)
@@
$\Pr\left(E_{2}^{c}\right)$ $\leq\Pr\left(\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}|\sigma(\tilde{\nu})-\sigma(\hat{\nu})|+\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu})-\sigma(\nu)|\geq\frac{c_{1}}{2}\right)$
$\leq\Pr\left(\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}|\sigma(\tilde{\nu})-\sigma(\hat{\nu})|\geq\frac{c_{1}}{4}\right)+\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu})-\sigma(\nu)|\geq\frac{c_{1}}{4}\right)$
$\lesssim a_{1,d}e^{-a_{2,d,k}nc_{1}/4}+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4}).$
The final inequality follows directly from Conditions 1 and 2
@@
$\Pr\left(E_{3}^{c}\right)\leq\Pr\left(2\left(\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\hat{\mu}_{u}|+c_{3}\right)>4c_{3}\right)=\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\hat{\mu}_{u}|>c_{3}\right)\lesssim\zeta_{n,d}(c_{3}).$
@@
$\lesssim\zeta_{n,d}(\frac{c_{1}}{2}),$
@@
First, note that $\mu$ and $\sigma$ may satisfy Conditions 1 and 2 for different parameters, e.g., different sequences $\zeta$. However, we can define a set of joint parameters $\zeta,\kappa_{n},\ldots$ for which the bounds in Conditions 1 and 2 hold for both $\mu$ and $\sigma$, which we will use throughout the proof. Let $b_{k,u}^{-}=\inf_{\tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)}\sigma(\tilde{\nu})$, $\hat{\alpha}_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}-\inf_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}$, $\alpha_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\mu_{u}-\inf_{u\in\mathbb{S}^{d-1}}\mu_{u}$.
An application of Lemma 5.3 with $\phi=O$, and subsequently taking $k=4\log(1/2\delta)/\epsilon$ yields:
$\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)$ $\geq 1-\Pr\left(E_{1}\right)$
$\hskip 30.0pt-\inf_{k\in[n]}\left[\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k+1)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}+G\left(\log\frac{1}{2\delta}-\frac{\epsilon k}{2}\right)\right]$ (B.7)
$\geq 1-\Pr\left(E_{1}\right)-\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},4\log(1/2\delta)/\epsilon+1)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}-\delta,$ (B.10)
@@
$\leq\Pr\left(\frac{2}{c_{1}}\left[\tau\tilde{S}_{n,k}\right]\geq\eta/2|E_{2}\right)+\Pr\left(\frac{2}{c_{1}}\left[(\sigma)+\tilde{S}_{n,k}(\mu)\right]\geq\eta/2|E_{2}\right)$
$=\Pr\left(\tilde{S}_{n,k}\geq c_{1}\eta/4\tau|E_{2}\right)+\Pr\left((\sigma)+\tilde{S}_{n,k}(\mu)\geq c_{1}\eta/4|E_{2}\right).$
@@
$\qquad\lesssim a_{1,d}\exp\left(-a_{2,d,k}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4}).$
@@
$\Pr\left(E_{5}^{c}\right)\lesssim a_{1,d}e^{-a_{2,d,k}nc_{1}/4}+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4}\wedge c_{3}).$
@@
$\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)$ $\geq 1-\Pr\left(E_{1}\right)-\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},4\log(1/2\delta)/\epsilon+1)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}-\delta$
$\gtrsim 1-\Pr\left(E_{5}^{c}\right)-\left(a_{1,d}\exp\left(-a_{2,d,k}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4})\right)-\delta$
$\gtrsim 1-(a_{1,d}e^{-a_{2,d,k}nc_{1}/4}+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4}\wedge c_{3}))$
$\qquad-\left(a_{1,d}\exp\left(-a_{2,d,k}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n,d}(\frac{c_{1}}{4})\right)-\delta$
$\gtrsim 1-\exp\left(-a_{2,d,k}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)-\kappa_{n}-\zeta_{n,d}(\frac{c_{1}}{4})-\delta\qed.$
@@
Furthermore, by assumption, $\Pr\left(\tilde{\theta}=\perp\right)=1-\mathbb{E}\lambda(\hat{\nu},\epsilon,\delta)\leq\gamma/3$. Thus, it suffices to find $t$ such that
@@
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)\\ \lesssim\inf_{\begin{subarray}{c}0<s<c_{1}\tau/2-c_{3}\\ 0<r<c_{1}\tau/2-s-c_{3}\end{subarray}}\Bigg{[}\exp\left(\frac{\epsilon(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\\ +\zeta_{n,d}\left(s\wedge c_{3}\wedge c_{1}/2\right)\Bigg{]}.$ (B.12)
@@
Next we bound $\zeta_{n,d}(s\wedge c_{3}\wedge c_{1}/2)\leq\gamma/3$. By assumption, $\gamma>3\zeta_{n,d}(c_{1}/2\wedge c_{3})$, and so we consider only $\zeta_{n,d}(s)$ for $s<c_{1}/2\wedge c_{3}$. To this end, using the fact that $\zeta_{n,d}$ are non-increasing, we have that
$\zeta_{n,d}(s)\leq\gamma/3\implies t\gtrsim\frac{c_{2}}{c_{1}}\zeta_{n,d}^{-1}(\gamma/3).$
This completes the proof.
Lemma B.2.
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)\\ \lesssim\inf_{\begin{subarray}{c}0<s<c_{1}\tau/2-c_{3}\\ 0<r<c_{1}\tau/2-s-c_{3}\end{subarray}}\Bigg{[}\exp\left(\epsilon\frac{(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\\ +\zeta_{n,d}\left(s\wedge c_{3}\wedge c_{1}/2\right)\Bigg{]}.$
Proof of Lemma B.2.
Condition 2 directly gives that $\Pr\left(E_{6,s}^{c}\right)\lesssim\zeta_{n,d}(s)$. Using this, the identities established in (B.2), (B.3), and (B.4), and noting that $c_{1}\leq c_{2}$, yields that for all $s\geq 0$, we have
$\Pr\left((E_{7}\cap E_{6,s})^{c}\right)\lesssim\zeta_{n,d}(s\wedge c_{3}\wedge c_{1}/2).$ (B.17)
$\geq\frac{t-c_{3}-s}{2c_{2}}.$ (B.18)
$\leq\frac{r+c_{3}+s}{c_{1}/2}.$ (B.19)
Therefore, taking $B_{r}(x)$ to be the ball of radius $r$ centered on $x$, then
Next, note that by definition
[…]
Similarly, an application of (B.19) gives
[…]
Combining these two bounds gives
Lemma B.3.
Proof.
We use the fact that
Let $\hat{\alpha}_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}-\inf_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}$.
Lemma B.4.
Proof.
In particular, taking $z=v\hat{\mu}_{v}$ for $v\in\mathbb{S}^{d-1}$,
@@
Let $\lambda_{t,q_{1},q_{2}}=\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ z\in(\xi_{q_{1},u}-t,\xi_{q_{2},u}+t)\end{subarray}}f_{\nu_{u}}(z)$ and $\hat{\xi}_{q,u}=F^{-1}_{\hat{\nu}_{u}}(q)$.
@@
We now consider the event $E_{9}$. Recall that the set $\mathcal{N}_{\epsilon}$ is an $\epsilon$-net on $\mathbb{S}^{d-1}$ if for all $u\in\mathbb{S}^{d-1}$, there exists $v\in\mathcal{N}_{\epsilon}$ such that $\lVert u-v\rVert\leq\epsilon$. Letting $\mathcal{N}_{\epsilon}$ be an arbitrary $\epsilon$-net on $\mathbb{S}^{d-1}$, it is straightforward (see, e.g., [25, Lemma 4.4.1]) to show that for any $i,j\in[n]$,
@@
Furthermore, there exists $\mathcal{N}_{1/4}$ such that $|\mathcal{N}_{1/4}|\leq 9^{d}$ [25, see Corollary 4.2.13]. Applying these two facts, in conjunction with two applications of subadditivity of probability measures, yields
@@
$\leq 9^{d}(n-2m+1)\sup_{\begin{subarray}{c}u\in\mathcal{N}_{1/4}\\ m\leq j<n-m\end{subarray}}\Pr\left((X^{\top}u)_{(j+1)}-(X^{\top}u)_{(j)}\geq\frac{3C}{4n}\right)$ (C.6)
@@
Let $U_{(j),u}=F_{\nu_{u}}((X^{\top}u)_{(j)})$. Now, recall that for any set of $n$ independent standard uniform random variables $V_{1},\ldots,V_{n}$, for any $j\in[n-1]$, it holds that $V_{(j+1)}-V_{(j)}\stackrel{{\scriptstyle\text{d}}}{{=}}V_{(1)}$ and $V_{(1)}\sim Beta(1,n)$ [see, e.g., 19]. These two facts, combined inequality (C.10), results in
$\sup_{\begin{subarray}{c}u\in\mathcal{N}_{1/4}\\ m\leq j<n-m\end{subarray}}\Pr\left(A_{3,u}|E_{8}\right)$ $\leq\sup_{\begin{subarray}{c}u\in\mathcal{N}_{1/4}\\ m\leq j<n-m\end{subarray}}\Pr\left(U_{(j+1),u}-U_{(j),u}\geq 3CM/4n\right)$ (C.15)

### v2

@@ Appendix A Algorithm
Appendix B Proofs from Section 3
We prove the main theorems with a series of lemmas. Let $\hat{\alpha}_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}-\inf_{u\in\mathbb{S}^{d-1}}\hat{\mu}_{u}$ and $\alpha_{\mu}=\sup_{u\in\mathbb{S}^{d-1}}\mu_{u}-\inf_{u\in\mathbb{S}^{d-1}}\mu_{u}$. The first lemma bounds the values of $\mu$ and $\sigma$ when computed on contaminated datasets.
@@ Appendix B Proofs from Section 3
    $\Pr\Big{(}\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}\sigma(\tilde{\nu})\leq\frac{c_{1}}{2}\Big{)}\lesssim a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4}).$ (B.1)
@@
    $\Pr\left(\hat{\alpha}_{\mu}>4c_{3}\right)\lesssim\zeta_{n}(c_{3}).$ (B.2)
@@
    $\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\geq 2c_{2}\right)\lesssim\zeta_{n}(c_{2}).$ (B.3)
@@
    $\Pr\left(\inf_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})\leq c_{1}/2\right)\lesssim\zeta_{n}(\frac{c_{1}}{2}).$ (B.4)
@@
$\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}\sigma(\hat{\nu}_{u})>2c_{2}\right)$ $\leq\Pr\left(\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}|\sigma(\tilde{\nu})-\sigma(\hat{\nu})|+\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu})-\sigma(\nu)|\geq\frac{c_{1}}{2}\right)$
$\leq\Pr\left(\sup_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ \tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)\end{subarray}}|\sigma(\tilde{\nu})-\sigma(\hat{\nu})|\geq\frac{c_{1}}{4}\right)$
$\hskip 100.00015pt+\Pr\left(\sup_{u\in\mathbb{S}^{d-1}}|\sigma(\hat{\nu})-\sigma(\nu)|\geq\frac{c_{1}}{4}\right)$
$\lesssim a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4}).$
The final inequality follows directly from Conditions 1 and 2.
@@
$\Pr\left(\hat{\alpha}_{\mu}\geq 4c_{3}\right)$ $\leq\Pr\left(2\left(\sup_{u\in\mathbb{S}^{d-1}}|\mu_{u}-\hat{\mu}_{u}|+c_{3}\right)>4c_{3}\right)\lesssim\zeta_{n}(c_{3}).$
@@
$\lesssim\zeta_{n}(\frac{c_{1}}{2}),$
@@
The next lemma bounds the local $k$-sensitivity of the outlyingness function. Let $b_{k,u}^{-}=\inf_{\tilde{\nu}\in\{\hat{\nu}_{u}\}\cup\widetilde{\mathcal{M}}_{1}(\hat{\nu}_{u},k)}\sigma(\tilde{\nu})$.
Lemma B.2.
Proof.
We use the fact that
The next lemma bounds the change in volume of $A_{\tau}$, when $\tau$ is slightly increased.
Lemma B.3.
Proof.
In particular, taking $z=v\hat{\mu}_{v}$ for $v\in\mathbb{S}^{d-1}$,
We can now prove Theorem 3.1.
@@
First, note that $\mu$ and $\sigma$ may satisfy Conditions 1 and 2 for different parameters, e.g., different sequences $\zeta_{n}$. However, we can define a set of joint parameters $\zeta_{n},\kappa_{n},\ldots$ for which the bounds in Conditions 1 and 2 hold for both $\mu$ and $\sigma$, which we will use throughout the proof.
An application of Lemma 5.3 with $\phi=O$, and subsequently taking $k=2$ yields:
$\mathbb{E}Z$ $\geq 1-\Pr\left(E_{1}\right)$
$\hskip 30.0pt-\inf_{k\in[n]}\left[\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},k+1)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}+G\left(\log\frac{1}{2\delta}-\frac{\epsilon k}{2}\right)\right]$
$\geq 1-\Pr\left(E_{1}\right)-\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},3)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}-\delta e^{\epsilon},$ (B.6)
@@
$\leq\Pr\left(\frac{2}{c_{1}}\left[\tau\tilde{S}_{n,k}\right]\geq\eta/2|E_{2}\right)+\Pr\left(\frac{2}{c_{1}}\left[\tilde{S}_{n,k}(\mu)\right]\geq\eta/2|E_{2}\right)$
$=\Pr\left(\tilde{S}_{n,k}\geq c_{1}\eta/4\tau|E_{2}\right)+\Pr\left(\tilde{S}_{n,k}(\mu)\geq c_{1}\eta/4|E_{2}\right).$
@@
$\qquad\lesssim a_{1}\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4}).$
@@
$\Pr\left(E_{5}^{c}\right)\lesssim a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4}\wedge c_{3}).$
@@
$\mathbb{E}Z$ $\geq 1-\Pr\left(E_{1}\right)-\Pr\Big{(}\sup_{\begin{subarray}{c}\tilde{\nu}\in\widetilde{\mathcal{M}}_{1}(\hat{\nu},3)\\ x\in A_{\tau+\eta,\hat{\nu}}\cup A_{\tau+\eta,\tilde{\nu}}\end{subarray}}|O(x,\hat{\nu})-O(x,\tilde{\nu})|>\eta\Big{)}-\delta$
$\gtrsim 1-\Pr\left(E_{5}^{c}\right)-\left(a_{1}\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4})\right)-\delta$
$\gtrsim 1-(a_{1}e^{-a_{2}nc_{1}/4}+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4}\wedge c_{3}))$
$\qquad-\left(a_{1}\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)+\kappa_{n}+\zeta_{n}(\frac{c_{1}}{4})\right)-\delta$
$\gtrsim 1-\exp\left(-a_{2}\frac{c_{1}n}{4}\left[1\wedge\frac{\eta}{\tau\vee 1}\right]\right)-\kappa_{n}-\zeta_{n}(\frac{c_{1}}{4})-\delta\qed.$
The proof of Theorem 3.2 relies on the following concentration inequality.
Lemma B.4.
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)\\ \lesssim\inf_{\begin{subarray}{c}0<s<c_{1}\tau/2-c_{3}\\ 0<r<c_{1}\tau/2-s-c_{3}\end{subarray}}\Bigg{[}\exp\left(\epsilon\frac{(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\\ +\zeta_{n}\left(s\wedge c_{3}\wedge c_{1}/2\right)\Bigg{]}.$
Proof of Lemma B.4.
Condition 2 directly gives that $\Pr\left(E_{6,s}^{c}\right)\lesssim\zeta_{n}(s)$. Using this, the identities established in (B.2), (B.3), and (B.4), and noting that $c_{1}\leq c_{2}$, yields that for all $s\geq 0$, we have
$\Pr\left((E_{7}\cap E_{6,s})^{c}\right)\lesssim\zeta_{n}(s\wedge c_{3}\wedge c_{1}/2).$ (B.8)
$\geq\frac{t-c_{3}-s}{2c_{2}}.$ (B.9)
$\leq\frac{r+c_{3}+s}{c_{1}/2}.$ (B.10)
Therefore, taking $B_{r}(x)$ to be the ball of radius $r$ centered on $x$, then
Next, note that by definition
[…]
Similarly, an application of (B.10) gives
[…]
Combining these two bounds gives
We now prove Theorem 3.2.
@@
Furthermore, by assumption, $\Pr\left(\tilde{\theta}=\perp\right)=1-\mathbb{E}Z\leq\gamma/3$. Thus, it suffices to find $t$ such that
@@
$\Pr\left(\lVert\tilde{\theta}-\theta\rVert\geq t\big{|}\tilde{\theta}\neq\perp\right)\\ \lesssim\inf_{\begin{subarray}{c}0<s<c_{1}\tau/2-c_{3}\\ 0<r<c_{1}\tau/2-s-c_{3}\end{subarray}}\Bigg{[}\exp\left(\frac{\epsilon(c_{1}/2+2c_{2})(c_{3}+s)}{4\eta c_{1}c_{2}}-\frac{\epsilon t}{8\eta c_{2}}+\frac{\epsilon r}{2c_{1}\eta}+\log\frac{(2c_{2}\tau+s+c_{3})^{d}-t^{d}}{r^{d}}\right)\\ +\zeta_{n}\left(s\wedge c_{3}\wedge c_{1}/2\right)\Bigg{]}.$ (B.12)
@@
Next we bound $\zeta_{n}(s\wedge c_{3}\wedge c_{1}/2)\leq\gamma/3$. By assumption, $\gamma>3\zeta_{n}(c_{1}/2\wedge c_{3})$, and so we consider only $\zeta_{n}(s)$ for $s<c_{1}/2\wedge c_{3}$. To this end, using the fact that $\zeta_{n}$ are non-increasing, we have that
$\zeta_{n}(s)\leq\gamma/3\implies t\gtrsim\frac{c_{2}}{c_{1}}\zeta_{n}^{-1}(\gamma/3).$
This completes the proof. ∎
@@
Before proving Theorems 3.4 and 3.5, we first prove two results concerning the projected quantiles and spacings. Let $\lambda_{t,q_{1},q_{2}}=\inf_{\begin{subarray}{c}u\in\mathbb{S}^{d-1}\\ z\in(\xi_{q_{1},u}-t,\xi_{q_{2},u}+t)\end{subarray}}f_{\nu_{u}}(z)$ and $\hat{\xi}_{q,u}=F^{-1}_{\hat{\nu}_{u}}(q)$.
@@
We now consider the event $E_{9}$. Recall that the set $\mathcal{N}_{\epsilon}$ is an $\epsilon$-net on $\mathbb{S}^{d-1}$ if for all $u\in\mathbb{S}^{d-1}$, there exists $v\in\mathcal{N}_{\epsilon}$ such that $\lVert u-v\rVert\leq\epsilon$. Letting $\mathcal{N}_{\epsilon}$ be an arbitrary $\epsilon$-net on $\mathbb{S}^{d-1}$, it is straightforward (see, e.g., [26, Lemma 4.4.1]) to show that for any $i,j\in[n]$,
@@
Furthermore, there exists $\mathcal{N}_{1/4}$ such that $|\mathcal{N}_{1/4}|\leq 9^{d}$ [26, see Corollary 4.2.13]. Applying these two facts, in conjunction with two applications of subadditivity of probability measures, yields
@@
$\leq 9^{d}(n-2m+1)\sup_{\begin{subarray}{c}u\in\mathcal{N}_{1/4}\\ m\leq j<n-m\end{subarray}}\Pr\left((X^{\top}u)_{(j+1)}-(X^{\top}u)_{(j)}\geq\frac{3C}{4n}\right)$
@@
Let $U_{(j),u}=F_{\nu_{u}}((X^{\top}u)_{(j)})$. Now, recall that for any set of $n$ independent standard uniform random variables $V_{1},\ldots,V_{n}$, for any $j\in[n-1]$, it holds that $V_{(j+1)}-V_{(j)}\stackrel{{\scriptstyle\text{d}}}{{=}}V_{(1)}$ and $V_{(1)}\sim Beta(1,n)$ [see, e.g., 20]. These two facts, combined inequality (C.5), results in
$\sup_{\begin{subarray}{c}u\in\mathcal{N}_{1/4}\\ m\leq j<n-m\end{subarray}}\Pr\left(A_{3,u}|E_{8}\right)$ $\leq\sup_{\begin{subarray}{c}u\in\mathcal{N}_{1/4}\\ m\leq j<n-m\end{subarray}}\Pr\left(U_{(j+1),u}-U_{(j),u}\geq 3CM/4n\right)$
@@
Proof of Theorem 3.4.
Note that:
$\sup_{\tilde{\nu}\in\tilde{\mathcal{M}}_{1}(\hat{\nu})}S_{n,k}(\mu)\leq\frac{\max_{0\leq i\leq k+1}X_{(n-\alpha n+1+k-i)}-X_{(\alpha n+1-i)}}{n-2\alpha n}.$
Therefore, applying Lemma C.2 we have that
$\Pr\left(S_{n,3}(\mu)\geq\eta\right)$ $\lesssim\max_{0\leq i\leq 4}\Pr\left(\sup_{u}\frac{(X^{\top}u)_{(n-\alpha n+4-i)}-(X^{\top}u)_{(\alpha n+1-i)}}{n-2\alpha n}\geq\eta\right)$
$\lesssim e^{-Mn^{2}(1-2\alpha)\eta/4+d+\log(n(1-2\alpha))}+e^{-nM^{2}r^{2}+d\log(cn/d)}.$
This proves the assertion for the trimmed mean. We now focus on the trimmed absolute deviations. Consider some $\tilde{\nu}\in\tilde{\mathcal{M}}_{1}(\hat{\nu},3)$ and let $Y_{1},\ldots,Y_{n}$ be the observations that correspond to $\tilde{\nu}$. For a particular $u\in\mathbb{S}^{d-1}$, let $A_{1,u},A_{2,u}\subset[n]$ be the subset of indices which are not trimmed from $\hat{\nu}_{u}$ and $\tilde{\nu}_{u}$, respectively. Now, first, it follows from above that with probability $1-e^{-Mn(1-2\alpha)\log n/4+d+\log(n(1-2\alpha))}-e^{-nM^{2}r^{2}+d\log(cn/d)}$, we have that $S_{n,3}(\mu)=\sup_{u,\tilde{\nu}}|\hat{\mu}_{u}-\tilde{\mu}_{u}|\leq\eta$. Therefore, we have that for all $u\in\mathbb{S}^{d-1}$ and $i\in[n]$, it holds that
$|Y_{i}^{\top}u-\tilde{\mu}_{u}|\leq\begin{cases}sign(Y_{i}^{\top}u-\hat{\mu}_{u})(Y_{i}^{\top}u-\hat{\mu}_{u}+\hat{\mu}_{u}-\tilde{\mu}_{u})&|Y_{i}^{\top}u-\hat{\mu}_{u}|\geq\eta\\ |Y_{i}^{\top}u-\hat{\mu}_{u}+\hat{\mu}_{u}-\tilde{\mu}_{u}|&|Y_{i}^{\top}u-\hat{\mu}_{u}|\leq\eta\end{cases}.$
Using this representation, we have that for $i\in A_{1,u}\cap A_{2,u}$, we have that$||X_{i}^{\top}u-\hat{\mu}_{u}|-|Y_{i}^{\top}u-\tilde{\mu}_{u}||\leq 2\eta.$ Applying this inequality yields that
$S_{n,3}(\sigma)$ $=\sup_{u\in\mathbb{S}^{d-1}}\left|\sum_{i\in A_{1,u}}\frac{|X_{i}^{\top}u-\hat{\mu}_{u}|}{n(1-2\alpha)}-\sum_{i\in A_{2,u}}\frac{|Y_{i}^{\top}u-\tilde{\mu}_{u}|}{n(1-2\alpha)}\right|$
$\leq 3\sup_{u\in\mathbb{S}^{d-1}}\sup_{i\in[3]}\left|\frac{|X_{i}^{\top}u-\hat{\mu}_{u}|-|Y_{i}^{\top}u-\hat{\mu}_{u}|}{n(1-2\alpha)}\right|+2\eta$
$\lesssim\sup_{u\in\mathbb{S}^{d-1}}\sup_{i\in[3]}\left|\frac{X_{i}^{\top}u-Y_{i}^{\top}u}{n(1-2\alpha)}\right|+\eta$
$\lesssim\eta,$
where the last inequality comes from the fact that $S_{n,3}(\mu)\leq\eta$. It follows that there exists a universal constant $C>0$ such that
$\Pr\left(S_{n,3}(\sigma)\geq\eta\right)\leq\Pr\left(S_{n,3}(\sigma)\geq\eta/C\right),$
and this completes the proof. ∎
The proof of Theorem 3.5 is a consequence of two lemmas, which we present as follows.